OpenStudy (anonymous):

How does this work? Please help. y = cos x. And then i have a chart of 30 degrees, 60 degrees, and 90 degrees.

OpenStudy (anonymous):



OpenStudy (anonymous):

\(\cos(60)=\frac{1}{2}\) from the picture and also \(\cos(30)=\frac{\sqrt{3}}{2}\)

OpenStudy (anonymous):

What does cos have to do with the equation? Because what im most confused about is the cos and not know what to put into my calculator.

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

cosine in a right triangle is "adjacent over hypotenuse" if you make the hypotenuse 1, it is just the "adjacent" side

OpenStudy (anonymous):

look at the unit circle on the last page of the attached cheat sheet find 30 degrees, then look at the coordinates on the circle the first coordinate is cosine, the second coordinate is sine

OpenStudy (anonymous):

Oh okay. This helps. Thank you.